# DeepCell: Multiview Representation Learning for Post-Mapping Netlists

### v1

@@ V-C Experiment Settings
Our model is equipped into ABC as a plug-in and integrated into the command ‘runeco’, which is an efficient SAT-based solution that won the first place in the 2017 ICCAD contest problem A and added the SAT-based exact pruning method [19]. Usually, the number of signals used in the patches generated by the ECO problem is not so much. In the subsequent experiments, the maximum number of candidate signals we selected is 1000. In [19], the quantified Boolean formula (QBF) method that can be used when the SAT-based feasibility calculation time exceeds the range is introduced. DeepCell, as a generic plug-in, can also be added to these processes if the method code is open source.
@@ V-D Main Results
We verify the effectiveness of our model in the 2017 ICCAD contest problem A [21]. These benchmarks are using ISCAS-85/89 [22], ITC-99 [23], IWLS-2005 [24], OpenCore and LGSynth-93 benchmarks and real-world ECO problems from industrial design. As shown in Table III, the sections labeled ‘gate(F)’ and ‘gate(G)’ correspond to the number of gates in the original circuit and the golden circuit, respectively. The ‘w/o DeepCell’ and ‘w/ DeepCell’ sections lists the effects of ‘runeco’ before and after inserting DeepCell. We limit the maximum running time to two hours. These timeout units are labeled as ‘-’. It should be denoted that the time reported in the table includes both model inference time and ECO tool runtime. All experiments are performed on a single core of 2.10 GHz Intel(R) Core(TM) i7-14700F CPU with 32GB memory.
From Table III, we can observe that that integrating our DeepCell into the ECO tool achieves significant improvements. Specifically, it reduces the average cost of patches by 6% and the average number of gates by 15%. Additionally, the required runtime decreases by 15%, with no units experiencing increased costs or gate counts. Notably, in the top-performing unit 16, our model achieves a remarkable 27% reduction in cost and a 40% reduction in gate count. These results are attributed to the ability of effectively capturing node relationships, enabling it to calculate the probability of each node for every target in multi-target units.
@@
TABLE IV: Performance Comparison between w/ DeepCell and w/o DeepCell on Find Feasible ECO Solution
@@
For four units that encountered timeout and large runtime (unit 6, 10, 11, 19), we conduct further analysis in Table IV. Data regarding the cost, support size, and runtime for the patches created for the first target of these units are collected. We increase the limit time of this part of the test to three and a half hours. Except for unit 19, the other three units are able to provide a viable solution more quickly, and the quality of these solutions are superior.
VI Conclusion
We introduced DeepCell, a multiview representation learning framework for PM netlists. By integrating a GNN-based PM encoder with a pretrained AIG encoder, and leveraging the novel Mask Circuit Modeling mechanism, DeepCell captures rich and generalizable embeddings of PM netlists. Our experimental results demonstrate the effectiveness of DeepCell in functional ECO, achieving significant reductions in patch generation costs, gate count, and runtime. Future work will explore its adaptability to other critical EDA tasks and extend its scalability to industrial-scale designs.

### v2

@@ V-C Experiment Settings
Our model is equipped into ABC [17] as a plug-in and integrated into the command ‘runeco’, which is an efficient SAT-based solution that won the first place in the 2017 ICCAD contest problem A [41] and added the SAT-based exact pruning method [16]. To the best of our knowledge, this is the latest known open-source approach for functional ECO. Our method only predicts driving relations to prune candidate signals, make it compatible with the other ECO engines [23, 24]. As a general plug-in, DeepCell can be seamlessly integrated into these frameworks if source code is accessible.
In the following experiments, DeepCell demonstrates its effectiveness by selecting 1,000 driving nodes based on predicted probabilities as candidate signals. The benchmarks are using ISCAS-85/89 [42], ITC-99 [43], IWLS-2005 [44], OpenCore and LGSynth-93 benchmarks and real-world ECO problems from industrial design. All experiments including model inference are performed on a single core of 2.10 GHz Intel(R) Core(TM) i7-14700F CPU with 32GB memory.
@@ V-D Main Results
As shown in TABLE VI, the columns labeled ‘gate(F)’ and ‘gate(G)’ correspond to the number of gates in the original circuit and the golden circuit, respectively. The ‘w/o DeepCell’ and ‘w/ DeepCell’ sections lists the effects of ‘runeco’ before and after inserting DeepCell. The ABC tool runtime is labeled as time(s), the model runtime is labeled as model(s), and the overall(s) column accumulates both. The column target represents the number of patch insert locations. We limit the maximum running time to two hours. These timeout units are labeled as ‘-’.
Overall, DeepCell reduces the average cost of patches by 2.77% and the average number of gates by 11.60%. Additionally, the required runtime decreases by 25.63%, with no units experiencing increased costs or gate counts. Notably, in the top-performing unit 16, our model achieves a remarkable 27% reduction in cost and a 40% reduction in gate count. These results are attributed to the ability of effectively capturing node relationships, enabling it to calculate the probability of each node for every target in multi-target units.
@@
TABLE V: Performance Comparison between w/ DeepCell and w/o DeepCell on Find Feasible ECO Solution
@@
For four units encountering timeout (unit 6, 10, 11, 19), we conduct further analysis in TABLE V. The results regarding the cost, support size (number of candidate signals used), and runtime for the first single patch insertion of these units are collected. We increase the limit time to three and a half hours (12,600s). Except for unit 19, the other three units are able to provide a viable solution more quickly, and the quality of these solutions are superior. To be specific, unit 6 is timeout on w/o DeepCell ECO engine, but can be efficiently solved in 538.06s with DeepCell.
TABLE VI: Functional ECO Performance Comparison between w/ DeepCell and w/o DeepCell on ICCAD’17 Contest Benchmarks
Circuit name Circuit information w/o DeepCell w/ DeepCell
PI PO gate(F) gate(G) tatget cost gate time(s) cost gate model(s) time(s) overall(s)
unit 1 3 2 6 6 1 4 1 0.37 4 1 0.06 0.37 0.43
unit 2 157 64 1,120 1,219 1 17 4 0.92 17 4 0.41 1.01 1.42
unit 3 411 128 2,074 1,929 1 80 3 0.46 80 3 0.38 0.40 0.78
unit 4 11 6 75 77 1 32 5 0.49 32 2 0.10 0.55 0.65
unit 5 450 282 24,357 21,056 2 47 30 46.30 47 29 7.36 15.39 22.75
unit 6 99 128 13,828 11,812 2 - - - - - 5.13 - -
unit 7 207 24 2,944 1,721 1 284 2 8.50 284 2 0.47 6.88 7.35
unit 8 179 64 2,513 3,337 1 78 4 4.10 78 3 0.63 3.17 3.80
unit 9 256 245 5,849 4,657 4 50 35 1.33 50 26 0.65 1.70 2.35
unit 10 32 129 1,581 1,956 2 - - - - - 0.73 - -
unit 11 48 50 2,057 2,160 8 - - - 2,312 746 0.71 6,975.30 6,976.01
unit 12 46 27 13,804 821 1 104 1 1.02 104 1 0.36 0.71 1.07
unit 13 25 39 369 426 1 3,467 9 1.50 3,467 9 0.12 1.33 1.45
unit 14 17 15 1,981 1,006 12 95 41 5.43 95 41 0.23 6.72 6.95
unit 15 198 14 1,886 2,262 1 191 11 1.93 191 11 0.37 0.73 1.10
unit 16 417 214 2,371 9,324 2 278 15 16.58 204 9 1.98 14.67 16.65
unit 17 136 31 2,910 2,052 8 434 79 5.04 434 63 0.43 6.01 6.44
unit 18 245 100 4,860 3,881 1 18 1 5.14 18 1 0.97 2.44 3.41
unit 19 99 128 13,349 10,787 4 - - - - - 4.02 - -
unit 20 1,874 7,105 30,876 34,002 4 136 6 0.60 120 5 7.80 1.12 8.92
Geomean 147.88 9.21 4.49 143.78 8.14 0.61 3.34 4.16
Reduction - - - 2.77% 11.60% - 25.63% 7.23%
  • 1
    Exclude the easy cases w/o DeepCell time less than 1s and the hard cases cannot be addressed.
TABLE VII: Technology Mapping Performance Comparison between w/ DeepCell and w/o DeepCell on EPFL Benchmark Suite
Circuit PI/PO Skywater130nm Nangate45nm
Origin AIG w/o DeepCell w/ DeepCell w/o DeepCell w/ DeepCell
nodes levels area(nm2) delay(ps) area(nm2) delay(ps) area(nm2) delay(ps) area(nm2) delay(ps)
adder 256/129 1,310 97 4,790.84 7,314.47 4,900.95 7,362.27 966.38 1,311.01 964.78 1,577.49
arbiter 256/129 4,589 14 19,991.67 1,641.73 16,934.99 1,930.43 3,777.73 446.96 3,448.16 481.82
bar 135/128 3,600 11 9,972.06 3,848.12 9,658.01 3,658.24 2,328.56 2,011.48 1,810.66 1,974.39
cavlc 10/11 600 10 1,870.54 1,313.52 1,875.55 1,164.15 437.04 337.64 391.82 318.92
ctrl 7/26 85 6 332.82 827.57 357.84 602.28 64.37 199.91 67.83 155.97
dec 8/256 304 3 1,203.65 642.70 1,203.65 643.65 252.97 194.92 247.91 152.91
div 128/128 58,798 1,621 227,563.25 211,175.92 208,626.33 242,915.17 45,573.25 50,162.67 41,685.66 52,871.67
i2c 147/142 1,013 8 3,624.73 1,672.64 3,434.54 1,547.64 653.30 433.73 676.97 392.02
int2float 11/7 208 9 758.23 731.89 670.64 730.83 151.09 180.92 137.52 209.27
log2 32/32 38,817 202 159,729.44 43,278.46 123,851.28 40,329.39 31,323.36 9,472.56 25,398.21 8,971.46
max 512/130 4,378 29 16,502.08 25,513.67 13,311.52 19,598.67 3,498.96 5,569.14 2,675.96 3,294.48
mem_ctrl 1204/1231 33,661 34 109,762.77 12,184.23 106,116.77 13,581.43 22,063.90 4,281.56 21,394.11 4,534.66
multiplier 128/128 31,817 125 135,399.86 17,563.55 102,020.34 18,984.09 26,249.41 4,059.93 20,924.09 4,120.63
priority 128/8 457 11 1,527.72 1,813.81 1,569.00 1,199.12 337.29 300.68 323.46 360.82
router 60/30 159 12 469.20 1,385.41 505.48 1,094.54 96.56 354.57 101.61 274.00
sin 24/25 6,896 100 32,406.08 20,786.12 23,842.87 18,327.04 6,750.55 4,593.38 4,831.36 3,775.06
sqrt 128/64 30,670 1,637 119,448.30 406,476.56 105,303.49 420,115.69 24,054.11 95,101.24 21,038.21 91,700.94
square 64/128 17,405 73 74,765.45 7,488.54 66,362.40 6,660.44 12,773.05 2,118.08 12,783.16 1,729.88
voter 1001/1 9,375 48 52,293.90 5,474.38 40,252.36 5,677.99 10,948.29 1,230.04 7,580.47 1,287.86
Geomean 11,676.95 6,336.12 10,586.68 5,939.16 2,366.03 1591.05 2115.02 1495.16
Area Delay Product (ADP) 73,986,525.76 62,875,983.61 3,764,465.11 3,162,290.41
Reduction - 15.02% - 16.00%
VI Downstream Task: Technology Mapping
VI-A Preliminary
Technology mapping is a crucial step in chip design flow, where an abstract representation of a circuit (typically in AIG) is mapped onto a specific technology library, consisting of standard cells. The goal of technology mapping is to find an optimal realization of the circuit using the available library cells, minimizing cost metrics such as area, delay, and power consumption, while ensuring that the functionality of the original design is preserved.
Cut-based technology mapping [45] with delay estimation is one of the most common approaches for achieving optimized results through dynamic programming [46]. In this process, nodes in the AIG are processed in topological order, with the objective of finding a delay-optimal match for each node in the target technology library. The delay values used to evaluate the quality of these matches are derived from a load-independent delay model, which assigns fixed delay values to gates based on their input delays and gain-based approximations. However, this approach often results in suboptimal outcomes because the delay estimation is inaccurate due to the absence of the actual capacitive load of the gates. Unfortunately, the load cannot be fully determined during the mapping process, as the detailed fan-out connections are not yet available. Consequently, predicting the fan-out connections to estimate capacitive load has the potential to enhance technology mapping.
VI-B Model Finetuning
Since DeepCell captures the correlation between PM netlist and AIGs, we finetune the AIG encoder within DeepCell framework to predict the cell type implementing PO of the pre-mapping AIG cone. As illustrated in Fig. 4, given the AIG $\mathcal{G}^{A}$ and its corresponding PM netlist, we first use the AIG encoder to encode $\mathcal{G}^{A}$ into node-level embedding vectors $\mathbf{H}^{A}$. These embedding vectors are then processed using a pooling function, and an MLP (multi-layer perceptron) is applied to predict the cell type of the PO. The cell type prediction loss $L_{ctp}$ is defined in Eq. (17), where $y^{\prime}$ represents prediction and $y$ denotes the ground truth cell type of the PO in the PM netlist.
To prepare the training samples, we randomly select 8,000 cells from various PM netlists and record their corresponding cell types as labels. For each selected cell, the corresponding cone from the PIs to the cell is extracted and mapped into the AIG format. The pretrained AIG encoder is then fine-tuned for 60 epochs over 0.5 hours using 8 Nvidia A800 GPUs (about 4 GPU hours). To further investigate the flexibility and versatility of DeepCell, we finetune it for two different mapping flows, utilizing two different technology libraries: Skywater130nm (which was part of the pretraining) and Nangate45nm (an out-of-distribution library). After finetuning, both models achieve cell type prediction accuracy of over 90%.
Figure 4: DeepCell finetuning for technology mapping
$\begin{split}\mathbf{H}^{A}&=\Phi^{A}(\mathcal{G}^{A})\\ y^{\prime}&=MLP(\text{Pooling}(\mathbf{H}^{A}))\\ L_{ctp}&=BCE(y^{\prime},y)\end{split}$ (17)
VI-C Experiment Settings
Our DeepCell framework is used to enhance the ‘&nf’ command in ABC [17], which is currently recognized as the state-of-the-art open-source technology mapping implementations. DeepCell predicts the cell type connectivity for each target node to generate more accurate delay estimation. We validate our approach using the EPFL combinational benchmark suite [47]. The input AIGs undergo preprocessing through the following optimized synthesis flow prior to technology mapping: ‘resyn2rs; compress2rs; &sopb’, iterating 6 times. We exclude the benchmark circuit ‘hyp’ since a known bug in ABC [48]. Our experiments are conducted on two open-source libraries: Skywater130nm and Nangate45nm, load into ABC through the command ‘read_lib -G 250’.
The finetuned DeepCell is capable of delivering predictions for nearly all benchmark circuits within 1 second. As technology mapping itself is a fast and one-time effort, temporal comparisons are of little consequence. The experimental environment maintains the same configurations as Sec. V.
VI-D Main Results
As shown in TABLE VII, the technology mapping engine with DeepCell (w/ DeepCell) outperforms the baseline (w/o DeepCell). First, by predicting the actual load for more accurate delay estimation, DeepCell improves delay performance compared to the baseline engine. For example, 11 out of 19 benchmark circuits in the Skywater130nm library show delay optimization. Mapping ‘priority’ using the Skywater130nm library with the baseline engine (w/o DeepCell) results in a delay of 1,813.81 ps, while DeepCell reduces this to 1,199.12 ps, a 33.89% reduction. Second, the w/ DeepCell mapping engine unexpectedly achieves better area results than the baseline (&nf). On average, the area is reduced by 9.3% in the Skywater130nm library. We attribute this improvement to our mapping approach, which modifies the critical path and creates more room for area optimization in the ‘&nf’ command. Third, although DeepCell was not pretrained on circuits mapped in the Nangate45nm library, it still predicts the cell type with over 90% accuracy after fine-tuning on this library. This demonstrates DeepCell’s flexibility and transferability across different technology libraries. As a result, DeepCell achieves an overall ADP reduction of 15.02% in the Skywater130nm library and 16.00% in the Nangate45nm library.
VII Conclusion
We introduce DeepCell, a multiview circuit representation learning framework. By integrating GNN-based encoders to learn AIG and PM representations and fusing them in a self-supervised manner, DeepCell captures richer and more generalizable embeddings. DeepCell framework introduces a novel AI-driven EDA methodology particular for the key cross-stage or cross-view applications, an area that has not been fully addressed by previous approaches. Our demonstration of DeepCell in functional ECO achieves significant reductions in patch generation costs, gate counts, and runtime. Additionally, DeepCell shows strong performance in technology mapping, providing high-quality solutions with reduced area and delay. We believe that DeepCell’s ability to fuse diverse circuit representations will unlock new possibilities for more efficient and accurate chip design.